A recent Gallop poll showed that 672 of 1019 adults nationwide think that marijuana should be legalized

Find the margin of error that corresponds to a 95% confidence interval. 
Answer #1

Part a)

Margin of Error = Z(α/2) √ ( (p*q) / n) = 0.0291

Part b)
p̂ = X / n = 672/1019 = 0.6595


p̂ ± Z(α/2) √( (p * q) / n)
0.6595 ± Z(0.05/2) √( (0.6595 * 0.3405) / 1019)
Critical value Z(α/2) = Z(0.05/2) = 1.96 ( From Z table )
Lower Limit = 0.6595 - Z(0.05) √( (0.6595 * 0.3405) / 1019) = 0.6304
upper Limit = 0.6595 + Z(0.05) √( (0.6595 * 0.3405) / 1019) = 0.6886
95% Confidence interval is ( 0.6304 , 0.6886 )
( 0.6304 < P < 0.6886 )